Find the amount of tax and the tax rate. Round to the nearest hundredth of a percent. Cost of item:$142 selling price: $161.15 Tax amount:$ ? Tax rate= ?%
Found 2 solutions by Solver92311, Theo:
Answer by Solver92311(821)



Tax Amount: Selling Price minus Cost of Item

Tax Rate: (Tax Amount divided by Cost of Item) times 100, rounded to two decimal places.

John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by Theo(13342)

you can't solve this problem for the tax or the tax rate because you need one to get the other.

the cost of the item is 142.
the selling price is 161.15
the profit is 161.15 - 142 = 19.15

you will be taxed on the profit of 19.15.

in order to find the tax, you need to know the tax rate.

in order to find the tax rate, you need to know the tax.
if you don't have one, you can't solve for the other.

the formula for tax is T = P * R

T is the tax
P is the profit
R is the tax rate

if the profit is 19.15, the formula becomes T = 19.15 * R
if you don't know R, then you can't find T.

if you take this formula and solve for R, it becomes R = T / 19.15
if you don't know T, then you can't find R.

if you don't know either, then you can't solve the problem.